OpenStudy (anonymous):

The price p (in dollars) and the quantity x sold of a certain product obey the demand equation: x = -20p + 500 where 0 < p ≤ 25 Revenue, R, can be found by the following equation: R = xp. Express R as a function of x hellppp?

OpenStudy (anonymous):

Well since you have to put R as a function of x, you'll have to get p by itself from x=-20p+500, once that you'll be able to plug it in for p in R=xp.

OpenStudy (anonymous):

I got R = -(1/20)x2 + 50x is this right:/

OpenStudy (anonymous):

Well let me put in in steps, x = -20p + 500 minus 500 to the other side x - 500 = -20p then divide by -20 to get p by itself (x-500)/-20 = p or p=\[-\frac{ x }{ 20 }+25\]Then put that expression in for p to get R = \[x(-\frac{ x }{ 20 }+25)\] Then distribute the x \[-\frac{ x ^{2} }{ 20 }+25x\] not sure where you got the 50x from. You were close though. :)

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

thanks